OpenStudy (anonymous):

MEDALS WILL BE GIVEN FOR THIS ONE EASY QUESTION 9) Solve the system using substitution. x = 2y + 6 x + y = 12

OpenStudy (anonymous):

2y+6 + y = 12

OpenStudy (arabpride):

first we have to set the equation to look like y = mx +b

OpenStudy (arabpride):

the equations***

OpenStudy (anonymous):

@DSS show your work ?

OpenStudy (anonymous):

I did. I substituted 2y +6 in for x in the 2nd equation.

OpenStudy (anonymous):

thank youuuu ( :

OpenStudy (anonymous):

What did you get

OpenStudy (anonymous):

x+ 2 ?

OpenStudy (anonymous):

2y+6 + y = 12 3y + 6 = 12 3y = 6 y = 2

OpenStudy (anonymous):

tht is right bettylove

OpenStudy (anonymous):

That is not right.

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

@DSS what about this Solve the system using elimination. 3x – 9y = 18 x + 2y = 1

OpenStudy (anonymous):

No. y = 2 and x = 10...Please check your work memek12

OpenStudy (anonymous):

2y+6 + y = 12 3y + 6 = 12 3y = 6 y = 2

OpenStudy (anonymous):

3x – 9y = 18 3x -9y = 18 3x -9y = 18 x + 2y = 1 -3(x+2y=1) -3x-6y = -3 -15y = 15 y = -1 To find x, substitute -1 in for y and solve.

OpenStudy (anonymous):

omggggg thank you so much at @DSS you were a big help like omg